OpenStudy (anonymous):

Simplify each expression. Radical 5 * Radical 20

OpenStudy (anonymous):

\[\sqrt{5} * \sqrt{20}\]

OpenStudy (anonymous):

@agent0smith ?? :)

OpenStudy (anonymous):

\[\sqrt{5\times 20}\] is a start

OpenStudy (anonymous):

then it should be fairly easy

OpenStudy (anonymous):

so \[\sqrt{60}\]

OpenStudy (anonymous):

so I mean Square root 100 so 10??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thankyou! :)